OpenStudy (anonymous):

One of the largest rectangular omelets ever cooked had a perimeter of 80 feet. The length was 10 feet more than twice its width. How many feet wide was the omelet?

OpenStudy (ash2326):

@Xylomixyee Perimeter of a rectangle is \[2\times(L+W)\] L=length W=width here L=10+2W do you understand this?

OpenStudy (anonymous):

The width equals 10?

OpenStudy (ash2326):

Correct :D

OpenStudy (anonymous):

How do you set it up?

OpenStudy (anonymous):